# SPECIAL RELATIVITY

### v9

@@ 15.5 How to Solve Problems Involving Many Trajectory Kicks
We shall now discuss the situation where there are $n$ arbitrary spaced kickers, all different from one another in terms of the rotation angle introduced. Let us consider how we may apply covariant particle tracking in this circumstance, and try to understand what is happening when we have for example an undulator downstream of the kicker setup. Formally, if one wants to calculate the radiation from the undulator one should take into account all kicks in the electron trajectory, from the generation of the electron. However, this situation is not surprising, if one analyzes the general expression for the radiation field from a single electron Eq.(LABEL:revwied). In fact, we should note that, in general, one needs to know the entire history of the electron from $t^{\prime}=-\infty$ to $t^{\prime}=\infty$ since the integration in Eq.(LABEL:revwied) is performed (at least formally) between these limits. However, this statement should be interpreted physically, depending on the situation under study: integration should in fact be performed from and up to times when the electron does not contribute to the field anymore.
@@
We should pointed out that it is the electrodynamics theory, which ultimately decides what part of the particle trajectory is important for calculating undulator radiation and what part can be neglected. The most important, general statement concerning the relevant part of the particle trajectory, is that it must be calculated according to the covariant method (if one wants to use the usual Maxwell’s equations).
@@
Let us consider the ultrarelativistic assumption $1/\gamma^{2}\ll 1$, which is verified for synchrotron radiation setups. In general, the introduction of a small parameter in any theory brings simplifications. The ultrarelativistic approximation implies a paraxial approximation and Eq.(LABEL:revwied) can be simplified to Eq.(38). Suppose that we take a situation in which the rotation angle of the first bending magnet upstream of the undulator is much larger than $1/\gamma$. In other words, we now consider an electron moving along a standard synchrotron radiation setup. The electron enters the setup via a bending magnet, passes through a straight section, an undulator, and another straight section. Finally, it leaves the setup via another bend. Note that, although the integration in Eq.(38) is performed from $\infty$- ∞ to $\infty$, the only (edge) part of the trajectory into the bending magnets contributing to the integral is of order of the radiation formation length $L_{f}$. Mathematically, it is reflected in the fact that $\Phi_{T}(z^{\prime})$ in Eq.(38) exhibits more and more rapid oscillations as $z^{\prime}$ becomes larger than the formation length. At the critical wavelength the formation length is simply of order of $R/\gamma$, $R$ being the radius of the bend. That simply corresponds to an orbiting angular interval $\Delta\theta\simeq 1/\gamma$. Typically, the critical wavelength of the radiation from a bending magnet in synchrotron radiation source is about 0.1 nm and the formation length in this case is only few millimeters.
@@
Figure 68: Standard undulator radiation setup. When the electron passes through a bending magnet there is the synchrotron radiation, washing out the (Frank- Ginzburg) fields of the fast moving charge. At the exit of the bending magnet we have ”naked” electron. There is a process of formation of the field-dressed electron within the formation length from the very beginning of the straight section downstream the bending magnet. The field-dressed ultrarelativistic electron has a visible transverse size of order a few microns for third generation synchrotron radiation sources.
@@
We start with the description of the field formation process along the straight section downstream of the switcher, based on the covariant approach. First of all we have to synchronize distant clocks within the lab frame. The synchronization procedure that follows is the usual Einstein synchronization procedure. It is assumed that in the Lorentz lab frame the electron proceeds following a rectilinear trajectory with velocity $v$. This assumption is used as initial condition. Then we can analyze situation downstream the switcher by using the usual Maxwell’s equations.
When one analyzes the process of ”field-dressed” electron formation from the viewpoint of the non covariant approach, one assumes the same initial conditions (rectilinear trajectory with velocity $v$) for the electron motion. Then one solves the electrodynamics problem of fields formation by using the usual Maxwell’s equations. We already mentioned that the type of clock synchronization which results in time coordinate $t$ in an electron trajectory $\vec{x}(t)$ is never discussed in accelerator physics. However, we know that the usual Maxwell’s equations are only valid in the Lorentz frame. The non covariant approach is obviously based on a definite synchronization assumption, but this is actually a hidden assumption. In our case of interest, within the lab frame the Lorentz coordinates are then automatically enforced. So one should not be surprised to find that in this simple case of rectilinear motion there is no difference between covariant and non covariant calculations of the initial conditions at the undulator entrance.
@@
Let us now see what happens with a weak dipole magnet (a kicker), which is installed in the straight section upstream of the undulator and is characterized by a small kick angle $(\gamma\theta_{k})^{2}\ll 1$. What do we expect for the undulator radiation? At first glance the situation is similar to the switcher setup and the electron trajectory is again divided into two parts: before and after the kicker. The most important difference, however, is that electrodynamics now dictates that both trajectories are important for the calculation of the undulator radiation. When the electron passes through the kicker there is no synchrotron radiation (to be more precise, in this case radiation is indistinguishable from the self-electromagnetic fields of the electron), washing out the virtual radiation fields like in the switcher case. We expect that an electron that passes through a kicker is still ”field-dressed”, but we have an electron whose fields has been perturbed, and now include information about the acceleration with respect to an inertial frame.
According to the conventional theory, as usual for Newtonian kinematics, the Galilean vectorial law of addition of velocities is actually used. Non-covariant particle dynamics shows that the direction of the electron trajectory changes after the kick, while its speed remains unvaried. In contrast, covariant particle tracking, which is based on the use of Lorentz coordinates, yields different results for the trajectory of the electron. The electron speed decreases from $v$ to $v(1-\theta^{2}_{k}/2)$. This result is at odds with the prediction from non-covariant particle tracking, because the corrected relativistic addition law for non-parallel velocities is used to calculate the electron trajectory.
@@ 15.6 Synchrotron Radiation in the Case of Particle Motion on a Helix
The presence of red shift in bending magnet radiation automatically implies the same problem for conventional cyclotron radiation theory. In the ultrarelativistic limit, there are well-known analytical formulas that describe the spectral and angular distribution of cyclotron radiation emitted by an electron moving in a constant magnetic field having a non-relativistic velocity component parallel to the field, and an ultrarelativistic velocity component perpendicular to it. According to the conventional approach, exactly as for the bending magnet case, the angular-spectral distribution of radiation is a function of the total velocity of the particle due, again, to the Doppler effect. In contrast, the covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion in the magnetic field direction. It should be note that cyclotron-synchrotron radiation emission is one of the most important processes in plasma physics and astrophysics and our corrections are important for a much wider part of physics than that of synchrotron sources.
@@ 15.6.1 Existing Theory
Let us discuss in some detail the relativistic cyclotron radiation. Here we shall only give some final results and discuss their relation with the conventional synchrotron radiation theory from bending magnet. In the case of an uniform translation motion with non-relativistic velocity along the magnetic field direction (Fig. 69), a widely accepted (in astrophysics) expression for the angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit is given by (2)
@@
where $K_{1/3}$ and $K_{2/3}$ are the modified Bessel functions, $\xi=1/\gamma$, $\psi=\chi-\alpha$, ($\chi$ is the angle between $\vec{v}$ and $\vec{B}$ and $\alpha$ that between $\vec{n}$ and $\vec{B}$); the angle $\psi$ is clearly the angular distance between the direction of the electron velocity $\vec{v}$ and the direction of observation $\vec{n}$. Here the $\omega_{c}$ is defined by $3eB\gamma^{2}/(2mc)$.
@@
Actually we have already discussed radiation from an ultrarelativistic electron on a helical orbit in the previous section. Equation Eq. (96) is the result we worked out above for the bending magnet radiation from a single electron with angular deflection with respect to nominal orbit. Eq. ( 100) does not look the same as Eq. (96). It will, however, if we now define the small deflection angle $\eta_{y}=\pi/2-\chi$ and the observation angle $\theta_{y}=\pi/2-\alpha$ (the observer is also assumed to be located in the vertical plane tangent to the trajectory i.e. $\theta_{x},\eta_{x}=0$). The integrals in Eq. (96) can be expressed in terms of the modified Bessel functions:
@@
Then, making the necessary variable changes, the formula reduces to Eq.( 100).
@@ 15.6.2 Methodology of Solving Problems Involving Boosts
The calculation leading to Eq. ( 100) is rather elaborate. It is therefore desirable to have an independent derivation. The simplest way of analyzing the radiation for an ultrarelativistic helical motion makes use of the theory of relativity and involves practically no calculations. The way for computing the radiation in the case of uniform translation is simple. One describes a complicated situation by finding a coordinate system where the analysis is already done (radiation in the case of circular motion) and transforms back to the old coordinate system. The reference system $S^{\prime}$ in which the electron moves in circular motion can be transformed to a reference system $S$ in which the electron proceeds following a helical trajectory. Eq. ( 100) holds, indeed, in the frame $S$ for a particle whose velocity is $(v_{x},v_{y},v_{z})=(v_{0}\sin\chi\sin\phi,v_{0}\cos\chi,v_{0}\sin\chi\cos\phi)$. The Lorentz transformation, which leads to the value $v_{y}=v_{0}\cos\chi$ for the $y$-component of the velocity yields $(v_{x},v_{y},v_{z})=(v^{\prime}\sin\phi^{\prime}/\gamma_{y},v_{y},v^{\prime}\cos\phi^{\prime}/\gamma_{y})$, where $\gamma_{y}=1/\sqrt{1-v^{2}_{y}/c^{2}}$, $v^{\prime}$ is the velocity of the electron in the frame $S^{\prime}$ and the phase angle $\phi^{\prime}=\phi$ is invariant. This means that, in order to end up in $S$ with a transverse (to the magnetic field direction) velocity $v_{\perp}=v_{0}\sin\chi$, one must start in $S^{\prime}$ with $v^{\prime}=\gamma_{y}v_{0}\sin\chi$. In the ultrarelativistic approximation $\gamma_{\perp}^{2}=1/(1-v^{2}_{\perp}/c^{2})\gg 1$, and one finds the simple result $v_{0}=v^{\prime}$, so that a Lorentz boost with non-relativistic velocity $v_{y}$ leads to a rotation of the particle velocity $\vec{v}_{0}$ of the angle $\eta=\pi/2-\chi\simeq v_{y}/c\ll 1$ (if angle $\eta$ is small and $v_{0}\simeq c$, we would write $\gamma_{y}\sin\chi\simeq 1$). If one transforms the radiation field for a particle in a circular motion in the system $S^{\prime}$, one obtains the result that the effect of a boost amounts to a rigid rotation of the angular-spectral distribution of the radiation emitted by the electron moving with velocity $v_{0}$ on a circle that is, once more, Eq. ( 100) (3) . The notation ”” used here to indicate the velocity in the transverse to the magnetic field direction should not be confused with the ”” referring to an acceleration of the electron in the transverse direction in the proceeding sections.
@@
It comes out quite naturally that the covariant way of analyzing the radiation for helical motion considered above is based on the Lorentz transformation. In other words, within the lab frame the Lorentz coordinates are automatically enforced. It assumed that in the Lorentz lab frame the electron proceeds following a helical trajectory with velocity $v_{0}$. This is employed as initial condition. In the ultrarelativistic approximation a Lorentz boost along the field direction with non relativistic velocity $v_{y}$ leads to the circular motion of the electron with the same velocity $v_{0}$. Thus the single boost along the field direction will leave the radiation properties unchanged.
@@
Let us discuss the synchrotron radiation from a single electron with a kick with respect to the nominal orbit in $(x,z)$-plane. In this case, the acceleration yields a particle velocity increment $v_{y}=v\theta_{k}$ parallel to the $y$-axis and $\Delta v_{z}=-v\theta^{2}_{k}/2$ parallel to the $z$-axis. The restriction to the second order provides an essential simplicity of the calculations. We can use a sequence of two commuting non-collinear Lorentz boosts in order to discuss the particle motion downstream the kicker.
@@
When the kick is introduced, the covariant particle tracking predicts a change of the initial condition at the entrance of the synchrotron radiation setup. If we will keep the Lorentz coordinate system in the lab frame downstream of the kicker, we will find that the covariant velocity on the helical orbit after the kick decreases from $v$ to $v-v\theta^{2}_{k}/2$ and the covariant way of analyzing the radiation for a helical motion with covariant velocity $v_{0}=v-v\theta^{2}_{k}/2$ considered above will leads to a red shift in the critical wavelength.
@@
It should be note that the single passive Lorentz boost to the reference system $S^{\prime}$ discussed above is only a mathematical trick. A good way to think of this transformation is to regard it as a result of transformation to new variables. According to this kinematic transformation, the electron (and the observer) motion relative to the fixed stars does not changes. An absolute (i.e. physically real) acceleration is always a dynamical process. We must conclude that when we accelerate with respect to the fixed stars an electron in the lab inertial frame upstream the uniform magnetic field, this acceleration is absolute and the information about this acceleration is included into the covariant trajectory.

### v10

@@ 19.5 How to Solve Problems Involving Many Trajectory Kicks
We shall now discuss the situation where there are $n$ arbitrarily spaced kickers, all different from one another in terms of the rotation angle introduced. Let us consider how we may apply covariant particle tracking in this circumstance, and try to understand what is happening when we have for example an undulator downstream of the kicker setup. Formally, if one wants to calculate the radiation from the undulator one should take into account all kicks in the electron trajectory, from the generation of the electron. However, this situation is not surprising, if one analyzes the general expression for the radiation field from a single electron Eq.(LABEL:revwied). We should note that, in general, one needs to know the entire history of the electron from $t^{\prime}=-\infty$ to $t^{\prime}=\infty$ since the integration in Eq.(LABEL:revwied) is performed (at least formally) between these limits. However, this statement should be interpreted physically, depending on the situation under study: integration should in fact be performed from and up to times when the electron does not contribute to the field anymore.
@@
We should point out that it is the electrodynamics theory, which ultimately decides what part of the particle trajectory is important for calculating undulator radiation and what part can be neglected. The most important, general statement concerning the relevant part of the particle trajectory, is that it must be calculated according to the covariant method (if one wants to use the usual Maxwell’s equations).
@@
Let us consider the ultrarelativistic assumption $1/\gamma^{2}\ll 1$, which is verified for synchrotron radiation setups. In general, the introduction of a small parameter in any theory brings simplifications. The ultrarelativistic approximation implies a paraxial approximation and Eq.(LABEL:revwied) can be simplified to Eq.(43). Suppose that we take a situation in which the rotation angle of the first bending magnet upstream of the undulator is much larger than $1/\gamma$. In other words, we now consider an electron moving along a standard synchrotron radiation setup. The electron enters the setup via a bending magnet and passes through a straight section, an undulator, and another straight section. Finally, it leaves the setup via another bend. Note that, although the integration in Eq.(43) is performed from $\infty$- ∞ to $\infty$, the only (edge) part of the trajectory into the bending magnets contributing to the integral is of the order of the radiation formation length $L_{f}$. Mathematically, it is reflected in the fact that $\Phi_{T}(z^{\prime})$ in Eq.(43) exhibits more and more rapid oscillations as $z^{\prime}$ becomes larger than the formation length. At the critical wavelength, the formation length is simply of the order of $R/\gamma$, $R$ is the radius of the bend. That corresponds to an orbiting angular interval $\Delta\theta\simeq 1/\gamma$. Typically, the critical wavelength of the radiation from a bending magnet in a synchrotron radiation source is about 0.1 nm. The formation length in this case is only a few millimeters.
@@
Figure 57: Standard undulator radiation setup. When the electron passes through a bending magnet there is synchrotron radiation, washing out the (Frank- Ginzburg) fields of the fast-moving charge. At the exit of the bending magnet, we have a ”naked” electron. There is a process of formation of the field-dressed electron within the formation length from the very beginning of the straight section downstream of the bending magnet. The field-dressed ultrarelativistic electron has a visible transverse size of the order of a few microns for third-generation synchrotron radiation sources.
@@
We start with the description of the field formation process along the straight section downstream of the switcher, based on the covariant approach. First of all, we have to synchronize distant clocks within the lab frame. The synchronization procedure that follows is the usual Einstein synchronization procedure. It is assumed that in the Lorentz lab frame the electron proceeds following a rectilinear trajectory with velocity $v$. This assumption is used as an initial condition. Then we can analyze the situation downstream of the switcher by using the usual Maxwell’s equations.
When one analyzes the process of ”field-dressed” electron formation from the viewpoint of the noncovariant approach, one assumes the same initial conditions (rectilinear trajectory with velocity $v$) for the electron motion. Then one solves the electrodynamics problem of field formation by using the usual Maxwell’s equations. We already mentioned that the type of clock synchronization that results in the time coordinate $t$ in an electron trajectory $\vec{x}(t)$ is never discussed in accelerator physics. However, we know that the usual Maxwell’s equations are only valid in the Lorentz frame. The noncovariant approach is obviously based on a definite synchronization assumption, but this is actually a hidden assumption. In our case of interest, within the lab frame, the Lorentz coordinates are then automatically enforced. So one should not be surprised to find that in this simple case of rectilinear motion, there is no difference between covariant and noncovariant calculations of the initial conditions at the undulator entrance.
@@
Let us now see what happens with a weak dipole magnet (a kicker), which is installed in the straight section upstream of the undulator and is characterized by a small kick angle $(\gamma\theta_{k})^{2}\ll 1$. What do we expect for the undulator radiation? At first glance, the situation is similar to the switcher setup, and the electron trajectory is again divided into two parts: before and after the kicker. The most important difference, however, is that electrodynamics now dictates that both trajectories are important for the calculation of the undulator radiation. When the electron passes through the kicker there is no synchrotron radiation (to be more precise, in this case, radiation is indistinguishable from the self-electromagnetic fields of the electron), washing out the virtual radiation fields like in the switcher case. We expect that an electron that passes through a kicker is still ”field-dressed”, but we have an electron whose field has been perturbed, and now include information about the acceleration with respect to an inertial frame.
According to the conventional theory, as usual for Newtonian kinematics, the Galilean vectorial law of addition of velocities is used. Non-covariant particle dynamics shows that the direction of the electron trajectory changes after the kick, while its speed remains unvaried. In contrast, covariant particle tracking, which is based on the use of Lorentz coordinates, yields different results for the trajectory of the electron. The electron speed decreases from $v$ to $v(1-\theta^{2}_{k}/2)$. This result is at odds with the prediction from noncovariant particle tracking because the corrected relativistic addition law for non-parallel velocities is used to calculate the electron trajectory.
@@ 19.6 Synchrotron Radiation in the Case of Particle Motion on a Helix
The presence of a redshift in bending magnet radiation automatically implies the same problem for conventional cyclotron radiation theory. In the ultrarelativistic limit, there are well-known analytical formulas that describe the spectral and angular distribution of cyclotron radiation emitted by an electron moving in a constant magnetic field having a non-relativistic velocity component parallel to the field, and an ultrarelativistic velocity component perpendicular to it. According to the conventional approach, exactly as for the bending magnet case, the angular-spectral distribution of radiation is a function of the total velocity of the particle due, again, to the Doppler effect. In contrast, the covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion in the magnetic field direction. It should be note that cyclotron-synchrotron radiation emission is one of the most important processes in plasma physics and astrophysics and our corrections are important for a much wider part of physics than that of synchrotron sources.
@@ 19.6.1 Existing Theory
Let us discuss in some detail the relativistic cyclotron radiation. Here we shall only give some final results and discuss their relation with the conventional synchrotron radiation theory from bending magnets. In the case of a uniform translation motion with non-relativistic velocity along the magnetic field direction (Fig. 58), a widely accepted (in astrophysics) expression for the angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit is given by  [^66] A widely accepted expression for the angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit were calculated in [69, 70]. At present, relativistic cyclotron radiation results are textbook examples (see e.g. [71]) and do not require a detailed description.
@@
where $K_{1/3}$ and $K_{2/3}$ are the modified Bessel functions, $\xi=1/\gamma$, $\psi=\chi-\alpha$, ($\chi$ is the angle between $\vec{v}$ and $\vec{B}$ and $\alpha$ that between $\vec{n}$ and $\vec{B}$); the angle $\psi$ is the angular distance between the direction of the electron velocity $\vec{v}$ and the direction of observation $\vec{n}$. Here the $\omega_{c}$ is defined by $3eB\gamma^{2}/(2mc)$.
@@
Actually, we have already discussed radiation from an ultrarelativistic electron on a helical orbit in the previous section. Equation Eq. (101) is the result we worked out above for the bending magnet radiation from a single electron with angular deflection with respect to the nominal orbit. Eq. ( 105) does not look the same as Eq. (101). It will, however, if we now define the small deflection angle $\eta_{y}=\pi/2-\chi$ and the observation angle $\theta_{y}=\pi/2-\alpha$ (the observer is also assumed to be located in the vertical plane tangent to the trajectory i.e. $\theta_{x},\eta_{x}=0$). The integrals in Eq. (101) can be expressed in terms of the modified Bessel functions:
@@
Then, making the necessary variable changes, the formula reduces to Eq.( 105).
@@ 19.6.2 Methodology of Solving Problems Involving Boosts
The calculation leading to Eq. ( 105) is rather elaborate. It is therefore desirable to have an independent derivation. The simplest way of analyzing the radiation for an ultrarelativistic helical motion makes use of the theory of relativity and involves practically no calculations. The way for computing the radiation in the case of uniform translation is simple. One describes a complicated situation by finding a coordinate system where the analysis is already done (radiation in the case of circular motion) and transforms back to the old coordinate system. The reference system $S^{\prime}$ in which the electron moves in circular motion can be transformed to a reference system $S$ in which the electron proceeds following a helical trajectory. Eq. ( 105) holds, indeed, in the frame $S$ for a particle whose velocity is $(v_{x},v_{y},v_{z})=(v_{0}\sin\chi\sin\phi,v_{0}\cos\chi,v_{0}\sin\chi\cos\phi)$. The Lorentz transformation, which leads to the value $v_{y}=v_{0}\cos\chi$ for the $y$-component of the velocity yields $(v_{x},v_{y},v_{z})=(v^{\prime}\sin\phi^{\prime}/\gamma_{y},v_{y},v^{\prime}\cos\phi^{\prime}/\gamma_{y})$, where $\gamma_{y}=1/\sqrt{1-v^{2}_{y}/c^{2}}$, $v^{\prime}$ is the velocity of the electron in the frame $S^{\prime}$ and the phase angle $\phi^{\prime}=\phi$ is invariant. This means that, in order to end up in $S$ with a transverse (to the magnetic field direction) velocity $v_{\perp}=v_{0}\sin\chi$, one must start in $S^{\prime}$ with $v^{\prime}=\gamma_{y}v_{0}\sin\chi$. In the ultrarelativistic approximation $\gamma_{\perp}^{2}=1/(1-v^{2}_{\perp}/c^{2})\gg 1$, and one finds the simple result $v_{0}=v^{\prime}$, so that a Lorentz boost with non-relativistic velocity $v_{y}$ leads to a rotation of the particle velocity $\vec{v}_{0}$ of the angle $\eta=\pi/2-\chi\simeq v_{y}/c\ll 1$ (if angle $\eta$ is small and $v_{0}\simeq c$, we would write $\gamma_{y}\sin\chi\simeq 1$). If one transforms the radiation field for a particle in a circular motion in the system $S^{\prime}$, one obtains the result that the effect of a boost amounts to a rigid rotation of the angular-spectral distribution of the radiation emitted by the electron moving with velocity $v_{0}$ on a circle that is, once more, Eq. ( 105)  [^67] The covariant way of analyzing the radiation for helical motion was considered in [72]. It is generally believed that $\vec{x}(t)=\vec{x}(t)_{cov}$ and this is the reason why in the [72] there is no distinction between the two (noncovariant and covariant) approaches to describe the electron motion on a helix downstream of the kicker setup.. The notation ”” used here to indicate the velocity in the transverse to the magnetic field direction should not be confused with the ”” referring to an acceleration of the electron in the transverse direction in the proceeding sections.
@@
It comes out quite naturally that the covariant way of analyzing the radiation for helical motion considered above is based on the Lorentz transformation. In other words, within the lab frame, the Lorentz coordinates are automatically enforced. It assumed that in the Lorentz lab frame the electron proceeds following a helical trajectory with velocity $v_{0}$. This is employed as an initial condition. In the ultrarelativistic approximation a Lorentz boost along the field direction with non relativistic velocity $v_{y}$ leads to the circular motion of the electron with the same velocity $v_{0}$. Thus the single boost along the field direction will leave the radiation properties unchanged.
@@
Let us discuss the synchrotron radiation from a single electron with a kick with respect to the nominal orbit in $(x,z)$-plane. In this case, the acceleration yields a particle velocity increment $v_{y}=v\theta_{k}$ parallel to the $y$-axis and $\Delta v_{z}=-v\theta^{2}_{k}/2$ parallel to the $z$-axis. The restriction to the second order provides an essential simplicity of the calculations. We can use a sequence of two commuting non-collinear Lorentz boosts in order to discuss the particle motion downstream of the kicker.
@@
When the kick is introduced, the covariant particle tracking predicts a change in the initial condition at the entrance of the synchrotron radiation setup. If we keep the Lorentz coordinate system in the lab frame downstream of the kicker, we will find that the covariant velocity on the helical orbit after the kick decreases from $v$ to $v-v\theta^{2}_{k}/2$ and the covariant way of analyzing the radiation for a helical motion with covariant velocity $v_{0}=v-v\theta^{2}_{k}/2$ considered above will lead to a red shift in the critical wavelength.
@@
It should be noted that the single passive Lorentz boost to the reference system $S^{\prime}$ discussed above is only a mathematical trick. A good way to think of this transformation is to regard it as a result of transformation to new variables. According to this kinematic transformation, the electron’s and the observer’s motion relative to the fixed stars does not change. An absolute (i.e. physically real) acceleration is always a dynamic process. We must conclude that when we accelerate with respect to the fixed stars an electron in the lab inertial frame upstream of the uniform magnetic field, this acceleration is absolute and the information about this acceleration is included in the covariant trajectory.